OpenStudy (anonymous):

Let θ be an angle in quadrant III such that =sinθ−14. Find the exact values of secθ and tanθ.

zepdrix (zepdrix):

\(\large\rm \theta=\sin\theta-14\) Did you make a typo somewhere?

OpenStudy (welshfella):

i think you must have..

OpenStudy (anonymous):

sin θ = - 1/4

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

ah thats more like it

OpenStudy (welshfella):



OpenStudy (welshfella):

sorry gtg right now

OpenStudy (anonymous):

Okay I got -4/-3.87 and -1/3.87

OpenStudy (anonymous):

Okay I got -4/-3.87 and -1/3.87

OpenStudy (anonymous):

sec = 4/-3.87 tan = -1/3.87

OpenStudy (anonymous):

@zepdrix

OpenStudy (welshfella):

the sec is correct the adjacent side = -3.87 so sec = -4/3.87

OpenStudy (welshfella):

but the tangent is positive because you are dividing 2 negative numbers

OpenStudy (welshfella):

tan = 1/3.87

OpenStudy (welshfella):